OpenStudy (zackattack350):

please help. for the given polynomial p(x) and the given c use the remainder theorem to find p(c) P(x)=8x^4-5x^2-2,1/3

OpenStudy (zackattack350):

can you help?

OpenStudy (zackattack350):

@DanJS

myininaya (myininaya):

is there difference between p and P?

myininaya (myininaya):

Also you don't need remainder theorem to find P(1/3)

OpenStudy (zackattack350):

thats how it was written...

myininaya (myininaya):

oh then how are p and P related?

OpenStudy (zackattack350):

i think its the same just a typo on the teachers part because its only in the written part once.

myininaya (myininaya):

P(1/3) can be found by replacing the x's with 1/3

OpenStudy (anonymous):

Evaluate P(1/3)

OpenStudy (campbell_st):

I think the remainder is 1/3 when you substitute x = c

OpenStudy (zackattack350):

so replace all x's with 1/3?

OpenStudy (anonymous):

\[P(\frac{ 1 }{ 3 }) = 8\left( \frac{ 1 }{ 3 } \right)^4 -5\left( \frac{ 1 }{ 3 } \right)^2-2\]

myininaya (myininaya):

is the question find P(c) given c is 1/3 or is the question find c such that P(c) is 1/3 ?*

OpenStudy (zackattack350):

use the remainder theorem to find P(C) so i believe its using 1/3 as a remainder

myininaya (myininaya):

if it is find c such that P(c) is 1/3 then @campbell_st is right

OpenStudy (campbell_st):

if I'm right, then its a tough question

myininaya (myininaya):

lol yep

OpenStudy (zackattack350):

i think its using the remainder

OpenStudy (anonymous):

c is given as 1/3 in the question. P(x) is given. You are asked to find P(c) using the remainder theorem. You may be overthinking it. The asker may just not have been consistent with using the upper case P.

OpenStudy (zackattack350):

they're the same. Its a typo on the question.

OpenStudy (zackattack350):

so do i just replace all X with 1/3?

OpenStudy (anonymous):

That's right.

myininaya (myininaya):

I don't see what this question has to do with remainder theorem like you can say use remainder theorem to find the remainder when P(x)/(x-c) by replacing the x in P(x) with the value for c... Like just evaluating P(c) alone is not using the remainder theorem

OpenStudy (lochana):

according to remainder theorem \[f(x) = (x-c)g(x) + a\] now you can easily find the solution 8x^4-5x^2-2 = (x-c)g(x) + a p(c) = 0 + 1/3 p(c) = 1/3

OpenStudy (anonymous):

The remainder theorem states that when dividing a polynomial P(x) by a binomial (x-c), the remainder is simply P(c). I believe that's what this question is asking. But I admit it's written in a very unusual way since there is no mention of the binomial x-c

myininaya (myininaya):

so maybe they want us to do synthetic division or long division

myininaya (myininaya):

\[\frac{P(x)}{x-\frac{1}{3}}\]

myininaya (myininaya):